# Chiral anomaly-induced nonlinear Hall effect in spin-orbit coupled noncentrosymmetric metals

### v2

@@ I Introduction
Weyl fermions, ever since it has been discovered in the context of high energy physics, took a long time to be realized in condensed matter systems as quasiparticles  [1, 2, 3, 4, 5]. Its realization as collective excitations in condensed matter systems, provides a realistic platform to test the prediction of the Weyl equation originally introduced in the high energy context [6, 7]. These materials are dubbed Weyl semimetals and are a very focused area of investigation among researchers due to their fascinating properties that were generally absent in other materials. Some of the very celebrated properties include open Fermi arcs  [2], anomalous Hall [8, 9] and Nernst effects [10, 11, 12], planar Hall and Nernst effects [13, 14], and most important one is the manifestation of chiral or Adler-Bell-Jackiw anomaly [15, 16, 17, 18, 19, 20, 20, 21, 22, 23, 24]. Weyl fermions have a very intrinsic property associated with them, i.e., chirality $\chi$ ($\pm 1$), which can be calculated by performing a closed surface integral of Berry flux around the band touching points (Weyl point). The number of Weyl Fermions associated with a particular chirality is fixed, leading to the electronic as well as chiral charge conservation. However, this does not hold when the system is brought under external gauge fields such as electric and magnetic fields. This is called chiral anomaly(CA).
@@
Weyl semimetals such as TaAs show negative magnetoresistance (positive magnetoconductance) induced by chiral anomaly [25, 26, 9] that has been verified experimentally [27, 23, 28]. In the absence of an electric field, an equivalent number of Weyl fermions exhibiting opposite chiralities move in opposite directions (aligned with the external magnetic field). This configuration leads to a null net charge current. However, when an electric field is applied parallel to the magnetic field, a discernible chemical potential difference arises between Weyl fermions with opposite chiralities, a consequence of the chiral anomaly [25]. This results in an imbalance between the two streams of Weyl fermions and, consequently, gives rise to a net charge current represented by $\mathbf{j}\propto(\mathbf{E}\cdot\mathbf{B)}\mathbf{B}$.
@@
Recently, CA was predicted to occur in materials where only one effective Weyl node is present with two disjointed and concentric Fermi surfaces having opposing Berry flux associated with it [29, 30, 31, 32]. These Weyl nodes are dubbed as ‘Kramers-Weyl nodes (KWNs) and are found at time-reversal-invariant momentum (TRIM) points [30, 31, 32]. The materials that hosts such nodes are called spin orbit coupled noncentrosymmetric metals (SOC-NCMs). Like WSMs, they too have a $\mathbf{k}\cdot\sigma$ term in the Hamiltonian but with an additional $\mathbf{k}^{2}$ term. However, unlike WSMs, the $\mathbf{\sigma}$ matrices here represents real spins. The Chern number turns out to be non-zero corresponding to the two Fermi surfaces, leading to a total sum still equal to zero [33, 34]. The recent findings advance the understanding of the CA induced linear as well as non-linear transport properties of SOC-NCMs [35, 36, 37, 38, 39, 40]. Although CA is demonstrated by both SOC-NCMs and WSMs, their transport responses are remarkably distinct [35, 36]. The non-linear Hall effect in WSM induced by chiral anomaly has been introduced by Rui-Hao Li et al. in Ref. [41]. Authors have found that the inversion symmetry-broken WSM having a tilt of the Weyl cone will have a nonlinear Hall effect induced by the chiral anomaly and dubbed it as ‘Chiral anomaly induced nonlinear Hall effect (CNLH)’. This CNLH is the combined effect of anomalous velocity $v^{\chi}_{a}=-\frac{e}{\hbar}\mathbf{E}\times\mathbf{\Omega^{\chi}_{k}}$ and chiral anomaly [42, 43]. To get CNLH, one has to integrate anomalous velocity on the Fermi surface [41, 36]. The CNLH effect is nonzero provided that the Fermi surface is asymmetric and the Hamiltonian has broken inversion symmetry. The tilt of the Weyl cone provides an asymmetric Fermi surface about the projection of the Weyl node on the Fermi surface [41]. Here, it is crucial to note that CNLH is distinct from CNLH caused by the Berry curvature dipole (BCD) [44, 45], as the latter can persist even without an external magnetic field.
@@
In previous works related to CNLH  [46, 47, 48], it is inherently assumed that the interband scattering rate is much lower than the intraband scattering rate. However, recent studies show that it is crucial to include interband scattering to naturally explain chiral anomaly induced charge transport [49, 50, 51, 10, 52, 11, 14, 53, 36, 35, 54]. Motivated by the impact of CA on transport properties in WSMs, under semiclassical treatment with chiral charge as well as number conservation, we have done a deeper analysis of CNLH in SOC-NCMs by going beyond the constant relaxation time approximation. In the present study, for the sake of completeness as well as for experimental usefulness, we have included magnetic as well as non-magnetic impurity scattering sites. The presence of magnetic impurity sites, which are magnetically active, can flip the spin of the incoming wave packet. This plays a very important role in the CA-assisted transport. In Fig. 2, we schematically draw the scattering mechanism for both cases: (i) chirality preserving (intraband) and (ii) chirality breaking (interband) scattering. Overlap of the wavefunction between incoming and outgoing wavefunctions has a different functional form: non-magnetic impurities ($\sigma_{0}$) preserve the spin structure, while magnetic impurities ($\sigma_{i},i=x,y,z$) cause spin-dependent scattering. This leads to interesting transport properties that provide a deeper understanding of the CNLH effect, that captures the missing facts in the previous studies, and will be more beneficial in elucidating the nonlinear conductivities that can appear via nonlinear transport measurement and nonlinear optical phenomena such as photocurrent and second harmonic production.
@@
For the case of inversion symmetry broken WSM (IWSMs) having a minimum of four nodes, one can get the imbalance of chiral charge density on nodes by adding a term in the Hamiltonian, as suggested in Ref. [41]. A pair of Weyl nodes having the same tilt orientation is sufficient to discuss the CNLH for IWSMs, as the properties for four nodes will be qualitatively similar and will be modified by the presence of two other nodes in magnitude. For WSMs, the CNLH is zero as long as the tilt of the Weyl cones is in opposite directions, as the individual contribution of each Weyl node gets canceled for opposite tilt orientation. For the case of SOC-NCMs, we interestingly get a non-zero CNLH even without the addition of anisotropy via tilt. This is due to the fact that the Fermi surfaces are of unequal size, and thus individual contributions for CNLH are unequal, the summation of which leads to a non-zero net CNLH even for the untilted case. The presence of OMM, tilt and spin Zeeman terms each enhanced the magnitude of CNLH. We first recover previous results: we have found that CNLH is quadratic in $B$, negative, symmetric about $B=0$, and does not show any sign reversal upon increment of interband scattering strengths. The presence of tilt in the dispersion introduced a linear in-$B$ term in the conductivity expressions, resulting in ‘weak and strong’ sign reversals. All four scattering sites taken show the same trends with respect to the tilt along the $x$-direction. We further notice that CNLH is highly anisotropic with respect to the direction of the tilt, in the presence of all four types of impurity sites, which has never been explored before. The article is arranged as follows. In section II, we discuss the mathematical formalisms used in the paper. In section II.1, we discussed the quasi-classical Boltzmann transport formalism(BTE) that incorporates charge conservation and momemtum dependent scattering. Section II.2 is devoted to the detailed study of the SOC-NCM Hamiltonian and how directional anisotropy in the presence of tilt is incorporated in the BTE. We discuss the main results of our paper in section III. Finally we conclude our discussions in section IV.
II Mathematical Modelling
II.1 Maxwell Boltzmann Transport formalism

### v3

@@ I Introduction
Weyl fermions, ever since it has been discovered in the context of high energy physics, took a long time to be realized in condensed matter systems as quasiparticles [bevan1997momentum, wan2011topological, Yan_2017, armitage2018weyl, burkov2011weyl]. Its realization as collective excitations in condensed matter systems, provides a realistic platform to test the prediction of the Weyl equation originally introduced in the high energy context [1929ZPhy...56..330W, abers2004quantum]. These materials are dubbed Weyl semimetals and are a very focused area of investigation among researchers due to their fascinating properties that were generally absent in other materials. Some of the very celebrated properties include open Fermi arcs [wan2011topological], anomalous Hall [yang2011quantum, burkov2014anomalous] and Nernst effects [sharma2016nernst, sharma2017nernst, liang2017anomalous], planar Hall and Nernst effects [nandy2017chiral, sharma2019transverse], and most important one is the manifestation of chiral or Adler-Bell-Jackiw anomaly [adler1969axial, nielsen1981no, nielsen1983adler, bell1969pcac, aji2012adler, zyuzin2012weyl, zyuzin2012weyl, son2012berry, goswami2015optical, fukushima2008chiral, goswami2013axionic]. Weyl fermions have a very intrinsic property associated with them, i.e., chirality $\chi$ ($\pm 1$), which can be calculated by performing a closed surface integral of Berry flux around the band touching points (Weyl point). The number of Weyl Fermions associated with a particular chirality is fixed, leading to the electronic as well as chiral charge conservation. However, this does not hold when the system is brought under external gauge fields such as electric and magnetic fields. This is called chiral anomaly(CA).
@@
Weyl semimetals such as TaAs show negative magnetoresistance (positive magnetoconductance) induced by chiral anomaly [son2013chiral, kim2014boltzmann, burkov2014anomalous] that has been verified experimentally [zhang2016linear, fukushima2008chiral, zyuzin2012topological]. In the absence of an electric field, an equivalent number of Weyl fermions exhibiting opposite chiralities move in opposite directions (aligned with the external magnetic field). This configuration leads to a null net charge current. However, when an electric field is applied parallel to the magnetic field, a discernible chemical potential difference arises between Weyl fermions with opposite chiralities, a consequence of the chiral anomaly [son2013chiral]. This results in an imbalance between the two streams of Weyl fermions and, consequently, gives rise to a net charge current represented by $\mathbf{j}\propto(\mathbf{E}\cdot\mathbf{B)}\mathbf{B}$.
@@
Recently, CA was predicted to occur in materials where only one effective Weyl node is present with two disjointed and concentric Fermi surfaces having opposing Berry flux associated with it [das2023chiral, bradlyn2016beyond, chang2018topological, hasan2021weyl]. These Weyl nodes are dubbed as ‘Kramers-Weyl nodes (KWNs) and are found at time-reversal-invariant momentum (TRIM) points [bradlyn2016beyond, chang2018topological, hasan2021weyl]. The materials that hosts such nodes are called spin orbit coupled noncentrosymmetric metals (SOC-NCMs). Like WSMs, they too have a $\mathbf{k}\cdot\sigma$ term in the Hamiltonian but with an additional $\mathbf{k}^{2}$ term. However, unlike WSMs, the $\mathbf{\sigma}$ matrices here represents real spins. The Chern number turns out to be non-zero corresponding to the two Fermi surfaces, leading to a total sum still equal to zero [cheon2022chiral, gao2022chiral]. The recent findings advance the understanding of the CA induced linear as well as non-linear transport properties of SOC-NCMs [varma2024magnetotransport, ahmad2025chiral, verma2019thermoelectric, sarkar2025symmetrydrivenintrinsicnonlinearpure, mandal2025disentangling, ahmad2025longitudinal]. Although CA is demonstrated by both SOC-NCMs and WSMs, their transport responses are remarkably distinct [varma2024magnetotransport, ahmad2025chiral]. The non-linear Hall effect in WSM induced by chiral anomaly has been introduced by Rui-Hao Li et al. in Ref. [PhysRevB.103.045105]. Authors have found that the inversion symmetry-broken WSM having a tilt of the Weyl cone will have a nonlinear Hall effect induced by the chiral anomaly and dubbed it as ‘Chiral anomaly induced nonlinear Hall effect (CNLH)’. This CNLH is the combined effect of anomalous velocity $v^{\chi}_{a}=-\frac{e}{\hbar}\mathbf{E}\times\mathbf{\Omega^{\chi}_{k}}$ and chiral anomaly [PhysRev.95.1154, xiao2010berry]. To get CNLH, one has to integrate anomalous velocity on the Fermi surface [PhysRevB.103.045105, ahmad2025chiral]. The CNLH effect is nonzero provided that the Fermi surface is asymmetric and the Hamiltonian has broken inversion symmetry. The tilt of the Weyl cone provides an asymmetric Fermi surface about the projection of the Weyl node on the Fermi surface [PhysRevB.103.045105]. Here, it is crucial to note that CNLH is distinct from other non-linear Hall effects caused by the Berry curvature dipole (BCD) [PhysRevB.97.201404, doi:10.1126/sciadv.1603266] or quadrupole [zhang2024third], as the latter can persist even without an external magnetic field.
@@
In previous works related to CNLH [li2021nonlinear, nandy2021chiral, zeng2022chiral], it is inherently assumed that the interband scattering rate is much lower than the intraband scattering rate. However, recent studies show that it is crucial to include interband scattering to naturally explain chiral anomaly induced charge transport [sharma2023decoupling, ahmad2021longitudinal, ahmad2023longitudinal, sharma2016nernst, sharma2017chiral, sharma2017nernst, sharma2019transverse, sharma2020sign, ahmad2025chiral, varma2024magnetotransport, ahmad2024geometry]. Motivated by the impact of CA on transport properties in WSMs, under semiclassical treatment with chiral charge as well as number conservation, we have done a comprehensive analysis of CNLH in SOC-NCMs by going beyond the constant relaxation time approximation. Additionally, we assume that the system is stationary and uniform so that the distribution function will merely be momentum–dependent. In the present study, for the sake of completeness as well as for experimental usefulness, we have included magnetic as well as non-magnetic impurity scattering sites. The presence of magnetic impurity sites, which are magnetically active, can flip the spin of the incoming wave packet. This plays a very important role in the CA-assisted transport. In Fig. 2, we schematically draw the scattering mechanism for both cases: (i) chirality preserving (intraband) and (ii) chirality breaking (interband) scattering. Overlap of the wavefunction between incoming and outgoing wavefunctions has a different functional form: non-magnetic impurities ($\sigma_{0}$) preserve the spin structure, while magnetic impurities ($\sigma_{i},i=x,y,z$) cause spin-dependent scattering. This leads to interesting transport properties that provide a deeper understanding of the CNLH effect, that captures the missing facts in the previous studies, and will be more beneficial in elucidating the nonlinear conductivities that can appear via nonlinear transport measurement and nonlinear optical phenomena such as photocurrent and second harmonic production.
@@
For the case of inversion symmetry broken WSM (IWSMs) having a minimum of four nodes, one can get the imbalance of chiral charge density on nodes by adding a term in the Hamiltonian, as suggested in Ref. [PhysRevB.103.045105]. A pair of Weyl nodes having the same tilt orientation is sufficient to discuss the CNLH for IWSMs, as the properties for four nodes will be qualitatively similar and will be modified by the presence of two other nodes in magnitude. For WSMs, the CNLH is zero as long as the tilt of the Weyl cones is in opposite directions, as the individual contribution of each Weyl node gets canceled for opposite tilt orientation. For the case of SOC-NCMs, we interestingly get a non-zero CNLH even without the addition of anisotropy via tilt. This is due to the fact that the Fermi surfaces are of unequal size, and thus individual contributions for CNLH are unequal, the summation of which leads to a non-zero net CNLH even for the untilted case, which is schematically presented in Fig. 1. The presence of orbital magnetic moment (OMM), tilt and spin Zeeman terms each enhanced the magnitude of CNLH. We first recover previous results: we have found that CNLH is quadratic in $B$, negative, symmetric about $B=0$, and does not show any sign reversal upon increment of interband scattering strengths. The presence of tilt in the dispersion introduced a linear in-$B$ term in the conductivity expressions, resulting in ‘weak and strong’ sign reversals. All four scattering sites taken show the same trends with respect to the tilt along the $x$-direction. We further notice that CNLH is highly anisotropic with respect to the direction of the tilt, in the presence of all four types of impurity sites, which has never been explored before. The article is arranged as follows. Section II is devoted to the detailed study of the SOC-NCM Hamiltonian and how directional anisotropy in the presence of tilt is incorporated in the BTE. We discuss the main results of our paper in section III. Finally we conclude our discussions in section IV. Remaining analytical calculations including discussion of the quasi-classical Boltzmann transport formalism(BTE) that incorporates charge conservation and momentum-dependent scattering for magnetic as well as non-magnetic scattering sites are presented in the Appendix.
II Spin-orbit coupled (SOC) noncentrosymmetric metals (NCMs) with weak Zeeman field
[…]
$+\hbar\vartheta(k_{x}t_{x}+k_{z}t_{z})\sigma_{0}.$ (2)
[…]
where, $t_{x}^{\prime}=\left(t_{x}+\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\cos{\gamma}}{(\hbar\vartheta)^{2}}\right)$, $t_{z}^{\prime}=\left(t_{z}+\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\sin{\gamma}}{(\hbar\vartheta)^{2}}\right)$, and the constant energy shift is given by $H_{0}=\left(\frac{\mathcal{M}}{\hbar\vartheta}\right)^{2}\frac{\hbar^{2}}{2m}+t_{z}\mathcal{M}\sin{\gamma}+t_{x}\mathcal{M}\cos{\gamma}$. Thus, under a coordinate transformation that shifts the origin, the Zeeman field effectively renormalizes the tilt components along the $x$- and $z$-directions as $t^{\mathrm{Z}}_{x}=\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\cos{\gamma}}{(\hbar\vartheta)^{2}}$ and $t^{\mathrm{Z}}_{z}=\frac{\hbar^{2}}{2m}\frac{2\mathcal{M}\sin{\gamma}}{(\hbar\vartheta)^{2}}$, respectively. Additionally, it introduces a tilt-dependent constant term to the Hamiltonian. This constant shift alters the Fermi surface profile as a function of the angle $\gamma$. Equation 3 retains the structural form of the Hamiltonian used for SOC-NCMs in Ref. [varma2024magnetotransport], allowing for straightforward diagonalization. The resulting energy eigenvalues are given by:
$+\hbar\vartheta(q_{x}t_{x}^{\prime}+q_{z}t_{z}^{\prime})+H_{\mathrm{0}},$ (4)
where $\chi=\pm 1$ denotes the two spin-orbit split bands. To obtain a constant energy contour—analogous to the case of Weyl semimetals—one must invert Eq. 4 with respect to $q$. In the transformed coordinate system, the expression for the orbital magnetic moment (OMM) retains a form similar to that of Weyl semimetals in the original coordinates. The inclusion of the OMM term renders Eq. 4 a cubic equation in $q$, yielding three possible solutions for each value of $\chi$. Among the three possible solutions, only one is physically acceptable, taking into consideration the validity of our low-energy model. The two other roots are not well behaved as $\mathbf{B}\to 0$. Moreover, the other “spurious” roots have a momentum range far from the low-energy window where our model does not hold. The physical solution for $q$ selected for analysis is plotted as a function of $\theta$ at $\phi=0$ in Fig. 3 for both $\chi=\pm 1$. The Fermi surface undergoes a chirality-dependent energy shift due to the OMM, i.e., $\epsilon^{\chi}_{\mathbf{k}}\rightarrow\epsilon_{\mathbf{k}}-\mathbf{m}^{\chi}_{\mathbf{k}}\cdot\mathbf{B}$. As a consequence of the modified dispersion relation, the velocity components are also altered and are given by the following expressions:
$v^{\chi}_{y}$ $=\left(\frac{\hbar}{m}\right)k_{y}+\vartheta\frac{k_{y}}{k}$
with $u^{\chi}_{2}=\frac{\chi e\vartheta B}{2\hbar}$.
III Results
Before discussing the CNLH, here we would like to briefly review the concept of different types of sign reversal. The sign of longitudinal magnetoconductivity (LMC) in Weyl materials has been the focus of extensive study in previous research. In the case of untilted Weyl semimetals, the chiral anomaly is expected to produce a positive LMC under conditions of weak interband scattering; however, with sufficiently strong interband scattering, the sign of LMC reverses [knoll2020negative, sharma2020sign, sharma2023decoupling]. Conversely, even a slight tilting of the Weyl cone can lead to negative LMC along a specific direction of the magnetic field, even in the regime of weak interband scattering. Despite the similarity in sign reversal, the underlying mechanisms in these scenarios are fundamentally distinct, giving rise to the classifications of ‘strong-sign-reversal’ and ‘weak-sign-reversal,’ as outlined in Ref. [ahmad2023longitudinal, varma2024magnetotransport]. We provide a brief review here. A general expression for the magnetoconductivity tensor can be expressed as [ahmad2023longitudinal]
$\sigma_{ij}(B)=\sigma_{ij}^{(0)}+(B-B_{0})^{2}\sigma_{ij}^{(2)},$ (6)
Figure 4: A schematic illustration depicting the characteristic regimes of magnetoconductivity $\sigma_{ij}$ in chiral Weyl semimetals, highlighting the phenomena of weak sign reversal, strong sign reversal, and the coexisting regime of strong-and-weak sign reversal, in contrast to the conventional quadratic-in-$B$ magnetoconductivity response [ahmad2023longitudinal, varma2024magnetotransport].
which integrates (i) normal quadratic $B$-dependence, (ii) linear-in-$B$ dependence with sign change along a specific magnetic field direction, and (iii) quadratic-in-$B$ dependence with a negative sign, within a unified framework. The features distinguishing ‘weak-sign-reversal’ include (i) $B_{0}\neq 0$, (ii) $\sigma_{ij}^{(0)}\neq\sigma_{ij}(B=0)$, and (iii) $\mathrm{sign}\;\sigma_{ij}^{(2)}>0$. In this situation, the vertex of the magnetoconductivity parabola shifts away from the origin, and the conductivity exhibits opposite signs for small positive and negative magnetic fields. However, the parabola’s orientation remains positive, i.e., $\mathrm{sign}\;(\sigma_{ij}^{(2)})>0$. In contrast, ‘strong-sign-reversal’ is identified by $\mathrm{sign}\;(\sigma_{ij}^{(2)})<0$, which signifies a complete inversion of the parabola’s orientation. Tilting of Weyl cones can result in ‘weak-sign-reversal,’ while strong interband scattering or strain typically leads to ‘strong-sign-reversal’ [ahmad2023longitudinal, varma2024magnetotransport]. Remaining possibility, i.e., completely reverse and shifted parabola results in ‘strong and weak-sign-reversal’. Fig. 4 illustrates the distinction between these three cases schematically.
Figure 5: The chiral anomaly-induced nonlinear Hall response (CNLH) is analyzed as a function of the interband scattering strength $\alpha$ and the angle $\gamma$, at a fixed magnetic field strength of $B=0.50\penalty 10000\ \mathrm{T}$. Panel (a) corresponds to the case of point-like non-magnetic impurity scattering, while panels (b), (c), and (d) represent magnetic impurity scattering with spin orientation aligned along the $\hat{x}-$, $\hat{y}-$, and $\hat{z}-$directions, respectively. The omission of the orbital magnetic moment (OMM) in our analysis affects only the overall amplitude—typically leading to a reduction—without altering the qualitative behavior. Hence, OMM has been included in all the present discussion.
III.1 CNLH in absence of tilt
In the absence of tilt in the band dispersion in SOC-NCMs, the chiral anomaly-induced nonlinear Hall (CNLH) conductance exhibits a quadratic dependence on the magnetic field $B$ and remains symmetric about $B=0$. In contrast to the linear-response regime—where a positive longitudinal magnetoconductance is typically observed due to the chiral anomaly—we consistently observe a negative magnetoconductance in the nonlinear regime. This behavior, characterized by a sign opposite to the linear response, is referred to here as a strong sign reversal. To examine the robustness of this feature, we analyze the behavior of CNLH conductance under varying strengths of interband scattering, denoted by $\alpha$. As expected, the magnitude of negative conductance decreases (i.e., becomes less negative) with increasing $\alpha$, indicating a suppression of anomaly-induced transport due to enhanced decoherence between bands. As established in earlier studies, the dependence of the conductance on the angle $\gamma$ between the magnetic field $\mathbf{B}$ and increasing $x$-direction shows that CNLH vanishes at $\gamma=0$ and $\gamma=\pi/2$, reaching a maximum at $\gamma=\pi/4$[ahmad2025chiral]. The $\gamma=\pi/2$ corresponds to the case when $\mathbf{E||B}$. The chiral anomaly source term is proportional to ($\mathbf{E}\cdot\mathbf{B}$), and is maximal for $\mathbf{E||B}$. However, the direction of current is controlled by the anomalous velocity $\mathbf{v_{a}}\propto(\mathbf{E}\times\Omega^{\chi})$, and by the remaining angular dependence of the integrand. For the untilted SOC-NCM (and in the absence of additional anisotropy in the scattering), the system has exact rotational symmetry in the $k_{x}-k_{y}$ plane when $\mathbf{B}$ is parallel to $\mathbf{E}$. The Berry curvature has a hedgehog-like structure, and the combination of $\mathbf{v_{a}}$, the distribution function, and the scattering kernel is such that the integrand for $J_{y}$ is odd under $\phi\to\phi+\pi$. As a result, the y-component of CNLH integrates to zero. Intuitively, when $\mathbf{E||B}$, the anomaly pumps charge between the Fermi surfaces along the field direction, but the induced transverse current (along the $y$-direction in our case) cancels pairwise due to the residual cylindrical symmetry. There is no preferred “Hall axis” in the plane perpendicular to $\mathbf{E}$, and hence the net CNLH $J_{y}$ must vanish. At $\gamma=0$, the projection $\mathbf{E}\cdot\mathbf{B}$ associated with the chiral anomaly becomes zero, thereby eliminating the anomaly-induced transport response. Furthermore, we investigate the influence of impurity type on the magnitude and profile of the CNLH conductance. We consider four distinct impurity scattering profiles: non-magnetic ($\sigma_{0}$) and magnetic impurities polarized along orthogonal directions ($\sigma_{x}$, $\sigma_{y}$, and $\sigma_{z}$). Although the overall qualitative behavior of the CNLH signal remains consistent across these impurity types, we observe significant quantitative variations. In particular, magnetic impurity types $\sigma_{x}$ and $\sigma_{y}$ lead to a larger magnitude of conductance, as illustrated in Fig. 5, whereas $\sigma_{0}$ and $\sigma_{z}$ impurities yield comparatively smaller conductance amplitudes.
Figure 6: The chiral anomaly-induced nonlinear Hall conductivity (CNLH) is analyzed as a function of the dimensionless interband scattering strength $(\alpha)$, for various values of the tilt parameter $t_{z}$, under a fixed magnetic field $B=0.50\penalty 10000\ \text{T}$ oriented along the direction $\gamma=\pi/4$. The arrows indicate the progression of $t_{z}$ from $0.98\,v_{F}$ to $0.98\,v_{F}$, illustrating the effect of tilt modulation on the nonlinear transport response.
Figure 7: The chiral anomaly-induced nonlinear Hall response (CNLH) is analyzed as a function of tilt along the $z$-axis ($t_{z}$) for various values of interband scattering strength ($\alpha$), under a fixed magnetic field $B=0.50\penalty 10000\ \text{T}$ oriented at an angle $\gamma=\pi/4$. Similar to the case with finite $t_{x}$, the CNLH vanishes at a finite, nonzero value of the tilt parameter $t_{z}$, indicating a symmetry-breaking threshold. For all four impurity scattering configurations considered—encompassing scalar and magnetic disorder profiles—the CNLH exhibits an approximately linear and monotonic dependence on $t_{z}$ within the parameter regime explored.
III.2 CNLH in the presence of tilt along the z-direction
III.3 CNLH in the presence of tilt along the x-direction
Figure 8: The chiral anomaly-induced nonlinear Hall conductivity (CNLH) is plotted as a function of the dimensionless interband scattering strength parameter ($\alpha$) for various values of the tilt parameter ($t_{x}$), under a fixed magnetic field $B=0.50\penalty 10000\ \text{T}$ applied at an oblique angle characterized by $\gamma=\pi/4$ with respect to $\hat{x}$. The dashed curves correspond to negative tilt configurations, i.e., $t_{x}/v_{F}<0$. Panels (a) and (d) exhibit qualitatively similar CNLH responses, whereas panels (b) and (c) reveal distinct deviations due to the interplay between the tilt-induced anisotropy and interband scattering.
Figure 9: The chiral anomaly-induced nonlinear Hall (CNLH) conductivity is plotted as a function of the tilt parameter ($t_{x}$) for various interband scattering strengths ($\alpha$), under a fixed magnetic field magnitude $B=0.50\penalty 10000\ \mathrm{T}$ oriented along the direction $\gamma=\pi/4$. Notably, the CNLH response vanishes at a finite (non-zero) value of tilt, indicating a nontrivial cancellation mechanism. With the exception of panel (b), the CNLH exhibits a predominantly linear and monotonic dependence on $t_{x}$, reflecting the symmetry-breaking role of tilt in modulating the anomalous transport.
IV Conclusion
Experimental proposal: There are several SOC-NCM candidates hosting Kramers-Weyl Fermions, that have sizable spin-orbit coupling and broken inversion symmetry that can potentially show CNLH as discussed in this paper. An exhaustive list of materials can be found in Refs. [wieder2022topological, chang2018topological]. From an experimental standpoint, our results suggest several clear signatures for detecting the chiral–anomaly–induced nonlinear Hall (CNLH) response in SOC–NCMs. An experimentally accessible indicator is the strong dependence of the CNLH magnitude on impurity type. Our analysis shows that transverse magnetic scatterers ($\sigma_{x}$, $\sigma_{y}$) substantially enhance the CNLH response, whereas nonmagnetic or longitudinal magnetic impurities ($\sigma_{0}$, $\sigma_{z}$) suppress it. This suggests that controlled magnetic alloying or implantation can serve as an effective handle to amplify the nonlinear signal and to distinguish the relative roles of intraband and interband scattering channels. We suggest an experimental setup as shown in Fig. 11. In order to experimentally probe CNLH effect, we propose the following: (i) apply an AC electric field along the z-axis and a static magnetic field in the xz-plane (as in our theory), and detect the second-harmonic Hall voltage (the CNLH signal) along the y-direction, (ii) rotate $\mathbf{B}$ and map out the angular dependence, looking for the predicted vanishing at $\gamma=0,\pi/2$ and the strong anisotropy with tilt, (iii) use controlled magnetic impurity doping or proximity to a magnetic material as shown in Fig. 11 to tune the relative strength of $\sigma_{0},\sigma_{x},\sigma_{y},\sigma_{z}$-type scattering channels, and observe the corresponding changes in the magnitude of CNLH, (iv) vary the carrier density (via chemical substitution) to shift the Fermi level and test the sensitivity of the effect to Fermi-surface asymmetry.The second harmonic Hall voltage refers to the Hall voltage measured at twice the driving frequency when the system is subjected to an ac electric field of the form $\bm{E}(t)=\bm{E}_{0}\cos(\omega t)$. Since the CNLH current is quadratic in the electric field, $J_{y}\propto E^{2}$, it naturally contains a component oscillating at frequency $2\omega$, which can be experimentally isolated using standard lock-in techniques. One can possibly ask as to how our proposed experimental set up can be used to probe a frequency dependent Hall voltage while our calculations assume a steady state with a definite CNLH conductivity. However, we wish to clarify that the assumption remains valid for the proposed ac measurement provided the driving frequency is low compared to the relevant scattering rates, i.e., $\omega\tau\ll 1$, where $\tau$ denotes the characteristic momentum and interband/intraband relaxation times. In this low-frequency regime, the system follows the external drive quasi-statically and relaxes continuously via impurity scattering, allowing a time-periodic steady state to be established.
Figure 11: Experimental setup to test the CNLH in SOC-NCMs having magnetic as well as non-magnetic scattering sites. The bottom layer represents the substrate while the middle layer is that of a magnetic material that controls the spin orientation of the magnetic scattering sites, represented by red in the top layers.
V Acknowledgments
Appendix A Maxwell Boltzmann Transport formalism